OpenStudy (anonymous):

The table shows the ratio between the number of books ordered and their cost: Number of Books Cost (dollars) 7 14 8 16 9 18 Find equivalent ratios when the number of books ordered is 1, 2, and 3. Write them as ordered pairs in the form (x-coordinate, y-coordinate). Starting from the origin, explain how to plot the three equivalent ratios on a coordinate grid. (10 points)

OpenStudy (poohdessey):

I don't understand the table .... can u upload the table

OpenStudy (anonymous):

here

OpenStudy (anonymous):

@hartnn peaseee help!

OpenStudy (anonymous):

@GIitter helpp

OpenStudy (anonymous):

@peak give it a try... pease...

OpenStudy (anonymous):

I rreally didn't get into ratios XD

OpenStudy (anonymous):

um im not all the way sure bout tht. but i can probably get someone who knows if u want

OpenStudy (anonymous):

kendricklamar2014 might be able to help.

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

@KarsonJamesNance @Ms-Brains

OpenStudy (anonymous):

Im sorry i dont understand it

OpenStudy (anonymous):

ok they are asking you to use the info you have with the 7 books = 14 dollars, 8 books = 16 dollars, and 9 books = 18 dollars. they want you to go back and find out what 1-6 books cost. so by subtracting 16 by 14 you will find how to find how many 1-6 books cost

MsBrains (ms-brains):

You're question is a mess, I m barely able to figure what you are asking. Sorry!

OpenStudy (anonymous):

hope i helped ^-^

OpenStudy (mathstudent55):

Each number of books and its cost can be written as a ratio. \(\dfrac{number~of~books}{cost} = \dfrac{7}{14} = \dfrac{8}{16} = \dfrac{9}{18}\) If you look at all those ratios written as fractions, they can be reduced. What does \(\dfrac{7}{14} \) reduce to?

OpenStudy (mathstudent55):

Once you reduce the ratio to simplest terms, then you can find how much one book costs. Then you'll be able to answer the question which is how much will 1, 2, and 3, books cost. Use this information as ordered pairs. For example, (7, 14), (8, 16), (9, 18).